OpenStudy (anonymous):

Solve: 4x – 9 = 15 1½ 6 8 24

OpenStudy (anonymous):

4x-9=15 add 9 to both sides 4x-9+9=15+9 4x=24 now divide by 4 x=6

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

and thanks for showing your work